As the 60-newton tension force acts upward and rightward on fido at an angle of 40 degrees find the components of the force
1
Expert's answer
2020-10-07T10:13:12-0400

"F_h=60\cdot\cos40°=45.96(N)"


"F_v=60\cdot\sin40°=38.6(N)"